OpenStudy (anonymous):

What is the sum of an 8–term geometric series if the first term is –11, the last term is 180,224, and the common ratio is –4?

OpenStudy (cwrw238):

S8 = -11 * ( 1 - (-4)^8) --------- 1 - (-4)

OpenStudy (cwrw238):

can u calculate that?

OpenStudy (anonymous):

144177

OpenStudy (cwrw238):

= 720885 / 5 = 144177